# Rethinking Machine Unlearning for Large Language Models

### v3

@@ 1 Introduction
Large language models (LLMs) have shown exceptional proficiency in generating text that closely resembles human-authored content. However, their ability to memorize extensive corpora may also lead to ethical and security concerns. These include societal biases and stereotyping (Bender et al., 2021; Motoki et al., 2023; Kotek et al., 2023), the generation of sensitive, private, harmful, or illegal content (Nasr et al., 2023; Wen et al., 2023; Karamolegkou et al., 2023; Patil et al., 2024), ease of jailbreaking (Wei et al., 2023; Zou et al., 2023; Liu et al., 2023a), and possible malicious use in developing cyberattacks or bioweapons (Barrett et al., 2023; Hendrycks et al., 2023). These concerns emphasize the need to adeptly and efficiently tailor pre-trained LLMs to suit diverse safety contexts while meeting specific requirements of users and sectors.
With the costly and prolonged training periods of LLMs, retraining these models to eliminate undesirable data effects is often impractical. Machine unlearning (MU) has emerged as an alternative to remove the influence of undesirable data and associated model capabilities from the pre-trained models (Cao & Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022; Si et al., 2023; Zhang et al., 2023a; Eldan & Russinovich, 2023; Yao et al., 2023). For example, MU is used as a strategy to prevent the generation of copyrighted material from the Harry Potter series in (Eldan & Russinovich, 2023). In the context of classification tasks, MU has been extensively studied by Ginart et al. (2019); Neel et al. (2021); Ullah et al. (2021); Sekhari et al. (2021); Golatkar et al. (2020); Jia et al. (2023). Yet, its application and understanding in LLMs remains limited, where models are typically used for generative tasks such as summarization, sentence completion, paraphrasing, and question answering. Therefore, this paper specifically concentrates on exploring the MU problems in LLMs, termed ‘LLM Unlearning’.
As data-model scales continue to grow, the emergence of LLM unlearning introduces new challenges and complexities, as will be elaborated on in Sec. 2. For example, current research efforts (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023c; Yao et al., 2023) suffer from a lack of standardized corpora and evaluation methods for LLM unlearning, often facing issues related to unlearning precision and efficiency. We also refer readers to Sec. 3 for a summary of existing LLM unlearning tasks.
@@
Although preliminary surveys of LLM unlearning have been provided by Si et al. (2023) and Zhang et al. (2023a), this paper is, to the best of our knowledge, the first to offer a thorough and in-depth analysis of LLM unlearning. Our contributions are summarized below. Fig. 1 provides an overview of the LLM unlearning landscape that we explore.
@@
(4) Forecasting: We offer insights into the future of LLM unlearning by identifying novel opportunities & challenges.
@@ 2 Preliminaries and Related Work
LLM unlearning has garnered attention for addressing trustworthiness concerns such as toxicity (Lu et al., 2022), copyright and privacy (Jang et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b), fairness (Yu et al., 2023), model editing (Ilharco et al., 2022; Zhang et al., 2023c), hallucination (Yao et al., 2023), and sensitive knowledge (Barrett et al., 2023; Hendrycks et al., 2023). In this section, we present a succinct overview of MU, tracing its journey from traditional ML models to the emerging challenges in LLMs.
MU for non-LLMs.
@@
The study of MU can be traced back to non-LLMs in response to data protection regulations such as ‘the right to be forgotten’ (Cao & Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2022). Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification (Ginart et al., 2019; Golatkar et al., 2020; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021), text-to-image generation (Gandikota et al., 2023; Zhang et al., 2023b; Kumari et al., 2023; Fan et al., 2024), federated learning (Liu et al., 2020; Wang et al., 2022; Che et al., 2023; Liu et al., 2023b; Halimi et al., 2022), and graph neural networks (Chen et al., 2022; Chien et al., 2022; Wu et al., 2023a).
@@
In the literature, ‘exact’ unlearning, which involves retraining the model from scratch after removing specific training data points, is often considered the gold standard. However, this approach comes with significant computational demands and requires access to the entire training set (Thudi et al., 2022). To address these challenges, many research efforts have shifted towards the development of scalable and effective approximate unlearning methods (Golatkar et al., 2020; Warnecke et al., 2021; Becker & Liebig, 2022; Thudi et al., 2022; Jia et al., 2023; Chen et al., 2023a). In addition, probabilistic methods with certain provable removal guarantees have been explored, often leveraging the concept of differential privacy (Ginart et al., 2019; Guo et al., 2019; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021).
Challenges of MU for LLMs.
LLM unlearning introduces new challenges and complexities.
First, LLMs are trained on massive amounts of data, which can unintentionally introduce biases and the memorization of personal and confidential information. Accordingly, it becomes challenging to precisely define and localize the ‘unlearning targets’, such as the subset of the training set or a knowledge concept that needs to be removed. Therefore, current studies on LLM unlearning (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023c; Yao et al., 2023) are typically context and task-dependent. There is a lack of standardized corpora for LLM unlearning.
Second, the growing size of LLMs and the rise of black-box access to LLM-as-a-service present challenges in developing scalable and adaptable MU techniques to LLMs (Bucknall & Trager, 2023; Casper et al., 2024a). This also affects performance evaluation, given the absence of retraining as a benchmark. To address these challenges, previous studies have proposed approaches like in-context unlearning (Pawelczyk et al., 2023) and fictitious unlearning (Maini et al., 2024), where the former enables unlearning on black-box models, and the latter provides alternatives to retraining.
Third, the scope of unlearning is often underspecified for LLMs. This issue is similar to challenges faced in model editing (Mitchell et al., 2022). For instance, effective unlearning should ensure that LLMs delete knowledge of the targeted data within the predefined scope while simultaneously maintaining its utility for data outside of this scope.
Fourth, despite the potential of LLM unlearning in diverse applications, there is a notable absence of comprehensive and reliable evaluation. For example, recent studies (Shi et al., 2023; Patil et al., 2024) have demonstrated that sensitive information can be reverse-engineered from an edited model, even if efforts were made to delete this information. This highlights the need for thorough and adversarial evaluations and the design of more mechanistic methods to guarantee the authenticity of unlearning.
@@ 3 Unpacking LLM Unlearning
In this section, we formalize the LLM unlearning problem and delve into its configuration. We will outline the inherent facts and complexities in its formulation.
Problem statement.
In light of the literature on unlearning (Bourtoule et al., 2021; Jia et al., 2023; Kurmanji et al., 2023), and its progression in LLMs (Pawelczyk et al., 2023; Yao et al., 2023; Ishibashi & Shimodaira, 2023; Maini et al., 2024), we define the problem of LLM unlearning below.
@@
Unlearning biographical knowledge
about fictitious authors.
@@
$\bullet$ Unlearning targets: Unlearning tasks may take on various forms and are closely related to the unlearning objectives. For instance, one could focus on data influence removal, while the other could emphasize model capability removal. Although these two aspects are intertwined, the former is often crucial for intellectual property (IP) protection, while the latter is more practical for AI alignment and safety. The literature identifies unlearning targets as specific data points, which could involve content containing harmful, unethical, or illegal language (Jang et al., 2022; Wu et al., 2023b). They have also been represented by higher-level unlearned knowledge, expressed through an unwanted text prompt or concept (Lu et al., 2022; Yao et al., 2023; Eldan & Russinovich, 2023). For example, Eldan & Russinovich (2023) defined the unlearning target as ‘Harry Potter’-related content, with the objective to avoid generating such content irrespective of where the content was learned: from the copyrighted material, blog posts, or news articles.
@@
$\bullet$ Influence erasure: Erasing the influence of unlearning targets and associated model capabilities requires a joint examination of both data and model influences rather than studies in isolation. Specifically, it is important to scrutinize the contributions of data sources to undesired model outputs, as well as the roles played by individual components within a model in generating these undesirable outcomes. This dual examination allows us to gain a more comprehensive understanding of the mechanisms driving these outputs, thereby facilitating the development of unlearning strategies to prevent them effectively. The objective of achieving complete influence erasure also implies the importance of robustness and generalization in unlearned behavior. When evaluating LLM unlearning, especially when using approximate methods shown in Table 1, a rigorous criterion is needed. Patil et al. (2024) underscore this viewpoint by demonstrating that sensitive information can persist in model weights after editing or unlearning, and by Maini et al. (2024) and Shi et al. (2023) who expose the limitations of conventional unlearning methods.
$\bullet$ Unlearning effectiveness: The effectiveness of LLM unlearning extends beyond merely diminishing the specific data influence or model capability. A crucial aspect of effectiveness is the concept of unlearning scope, as inspired by the editing scope (Mitchell et al., 2022). This has implications for the development of metrics to evaluate the effectiveness of unlearning and the design of effective unlearning methods. The unlearning scope defines the success of influence erasure by its capacity to accurately modify the model’s behavior for in-scope examples, which are either directly or indirectly related to unlearning targets, such as paraphrased text prompts to be degenerated (Zhong et al., 2023; Patil et al., 2024). At the same time, it must also ensure the generation consistency for out-of-scope examples, necessitating the preservation of model generation capabilities post-unlearning. Differentiating between in-scope and out-of-scope examples for unlearning is its own unsolved problem, as it requires determining when facts logically imply one another (Hase et al., 2023b; Cohen et al., 2023).
@@
$\bullet$ Unlearning efficiency & feasibility: Current research efforts focused on the development of rapid unlearning methods for pre-trained LLMs due to the significant re-training cost (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023). However, beyond computational efficiency, LLMs present additional efficiency challenges. These include the complexity and, at times, the infeasibility of pinpointing and attributing training data points designated for unlearning. Additionally, there is the challenge of executing unlearning in the context of black-box LLMs (Pawelczyk et al., 2023), where interactions with models are constrained to input-output queries.
Distinguishing LLM unlearning from conventional MU.
@@
Mathematical modeling.
Building upon the high-level formulation of LLM unlearning problems presented earlier, we next provide mathematical modeling details and discuss the associated design choices. To facilitate comprehension, we provide a commonly-used formulation of LLM unlearning problems below. While this may not be the sole or optimal approach to LLM unlearning, it incorporates several key elements that are essential to the problem setup.
@@
In the dataset setup of LLM unlearning, we typically assume access to a forget set ($\mathcal{D}_{\mathrm{f}}$), the influence of which should be eliminated in LLM generation. For instance, $\mathcal{D}_{\mathrm{f}}$ might consist of a collection of harmful or toxic prompt-response pairs designated for degeneration (Yao et al., 2023). Moreover, if the original training set is available, then $\mathcal{D}_{\mathrm{f}}$ can be composed of a subset of training data points related to the unlearning target. Alternatively, it can be generated using synthesized data points based on a higher-level unlearned knowledge concept, or it can be derived from a set of extracted training data points reverse-engineered from the given LLM itself. In practice, the forget set $\mathcal{D}_{\mathrm{f}}$ is not required to belong precisely to the LLM’s training corpus. Given the diversity and extensive size of the model’s training data, the content we aim to unlearn is more likely to represent a general concept. Thus, LLM unlearning needs to not only unlearn specific training samples but also generalize to similar samples that share common characteristics.
@@
Besides the forget set $\mathcal{D}_{\mathrm{f}}$, sometimes there is a need for a retain set ($\mathcal{D}_{\mathrm{r}}$), which contains samples that are not subject to unlearning and used to preserve the utility of the unlearned model. Through the lens of the unlearning scope we discussed earlier, the forget set ($\mathcal{D}_{\mathrm{f}}$) provides in-scope examples earmarked for unlearning, while the retain set ($\mathcal{D}_{\mathrm{r}}$) involves examples out of the unlearning scope. Some recent studies have also attempted to develop LLM unlearning approaches that operate independently of access to forget and/or retain sets (Pawelczyk et al., 2023; Li et al., 2023b).
@@
We next introduce the model and optimization setups for LLM unlearning. Unlearning is performed at the post-model training phase. As shown in (2), a common unlearning objective is to efficiently update the original pre-trained model so that the updated model can unlearn on $\mathcal{D}_{\mathrm{f}}$ while retaining its generation capability on $\mathcal{D}_{\mathrm{r}}$. In addition, another design element is unlearning response ($y_{\mathrm{f}}$), referred to as the response of an unlearned model to in-scope examples. For example, in the stateful LLM unlearning method aimed at erasing information related to ‘Who’s Harry Potter?’ (Eldan & Russinovich, 2023), the unlearning response is based on word replacements using generic translations, like substituting ‘Quidditch’ with ‘Skyball’, as part of the unlearning process. However, this type of approach may inadvertently blur the line between LLM hallucination and legitimate responses, highlighting the need for improvements in unlearning response design. Another choice is empty response (Wu et al., 2023b; Patil et al., 2024), given by the reject ‘I don’t know’ (Patil et al., 2024) or the customized response by ‘masking’ the unlearning information using specialized tokens such as $\blacksquare$’ in (Wu et al., 2023b). However, we need to ensure that the empty response targets only examples within the unlearning scope. Otherwise, frequent rejections may occur, potentially diminishing the user experience with LLMs.
4 LLM Unlearning Methods Revisited: Prior Art and Overlooked Principles
In this section, we begin with a review of existing MU (machine unlearning) methodologies adapted for LLM unlearning. Following that, we delve into the design factors that are often overlooked in the literature, including the intertwined data-model influence, the relationship with model editing, and adversarial training.
Review of existing unlearning principles.

### v4

@@ 1 Introduction
Large language models (LLMs) have shown exceptional proficiency in generating text that closely resembles human-authored content. However, their ability to memorize extensive corpora may also lead to ethical and security concerns. These include societal biases and stereotyping Bender et al. (2021); Motoki et al. (2023); Kotek et al. (2023), the generation of sensitive, private, harmful, or illegal content Nasr et al. (2023); Wen et al. (2023); Karamolegkou et al. (2023); Patil et al. (2024), ease of jailbreaking Wei et al. (2023); Zou et al. (2023); Liu et al. (2023b), and possible malicious use in developing cyberattacks or bioweapons Barrett et al. (2023); Hendrycks et al. (2023); Li et al. (2024a). These concerns emphasize the need to adeptly and efficiently tailor pre-trained LLMs to suit diverse safety contexts while meeting specific requirements of users and sectors.
With the costly and prolonged training periods of LLMs, retraining these models to eliminate undesirable data effects is often impractical Brown et al. (2020); Yao et al. (2024). Machine unlearning (MU) has emerged as an alternative to remove the influence of undesirable data and associated model capabilities from the pre-trained models Cao & Yang (2015); Bourtoule et al. (2021); Nguyen et al. (2022); Si et al. (2023); Zhang et al. (2023a); Eldan & Russinovich (2023); Yao et al. (2023). For example, MU is used as a strategy to prevent the generation of copyrighted material from the Harry Potter series Eldan & Russinovich (2023). In the context of classification tasks, MU has been extensively studied Ginart et al. (2019); Neel et al. (2021); Ullah et al. (2021); Sekhari et al. (2021); Golatkar et al. (2020); Jia et al. (2023). Yet, its application and understanding in LLMs remains limited, where models are typically used for generative tasks such as summarization, sentence completion, paraphrasing, and question answering. Therefore, this paper specifically concentrates on exploring the MU problems in LLMs, termed ‘LLM Unlearning’.
As data-model scales continue to grow, the emergence of LLM unlearning introduces new challenges and complexities, as will be elaborated on in Sec. 2. For example, current research efforts Lu et al. (2022); Jang et al. (2022); Ilharco et al. (2022); Eldan & Russinovich (2023); Wu et al. (2023b); Yu et al. (2023); Zhang et al. (2023c); Yao et al. (2023) suffer from a lack of standardized corpora and principled evaluation for LLM unlearning. Although preliminary surveys of LLM unlearning have been provided in Si et al. (2023); Zhang et al. (2023a), this paper is, to the best of our knowledge, the first to offer a thorough and in-depth review of LLM unlearning. The key contributions are summarized below. Fig. 1 provides an overview of the LLM unlearning landscape that we explore.
@@
(4) Forecasting: We offer insights into the future of LLM unlearning by identifying novel prospects and opportunities.
@@ 2 Related Work
LLM unlearning has garnered attention for addressing trustworthiness concerns such as toxicity Lu et al. (2022), copyright and privacy Jang et al. (2022); Eldan & Russinovich (2023); Wu et al. (2023b), fairness Yu et al. (2023), hallucination Yao et al. (2023), malicious usage Li et al. (2024a), and sensitive knowledge Barrett et al. (2023); Hendrycks et al. (2023). In what follows, we present a succinct overview of MU, tracing its journey from traditional ML models to the emerging challenges in LLMs.
MU for non-LLMs
@@
The study of MU can be traced back to non-LLMs in response to data protection regulations such as ‘the right to be forgotten’ Cao & Yang (2015); Hoofnagle et al. (2019); Bourtoule et al. (2021); Nguyen et al. (2022). Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification Ginart et al. (2019); Golatkar et al. (2020); Neel et al. (2021); Ullah et al. (2021); Sekhari et al. (2021), text-to-image generation Gandikota et al. (2023); Zhang et al. (2023b); Kumari et al. (2023); Fan et al. (2024b), federated learning Liu et al. (2020); Wang et al. (2022); Che et al. (2023); Liu et al. (2023c); Halimi et al. (2022), graph neural networks Chen et al. (2022b); Chien et al. (2022); Wu et al. (2023a), and recommendation Sachdeva et al. (2024); Chen et al. (2022a); Xu et al. (2023); Li et al. (2024b); Wang et al. (2024b).
@@
In the literature, ‘exact’ unlearning, which involves retraining the model from scratch after removing specific training data points, is often considered the gold standard. However, this approach comes with significant computational demands and requires access to the entire training set Thudi et al. (2022a). To address these challenges, many research efforts have shifted towards the development of scalable and effective approximate unlearning methods Golatkar et al. (2020); Warnecke et al. (2021); Becker & Liebig (2022); Thudi et al. (2022a); Jia et al. (2023); Chen et al. (2023a). In addition, probabilistic methods with certain provable removal guarantees have been explored, often leveraging the concept of differential privacy Ginart et al. (2019); Guo et al. (2019); Neel et al. (2021); Ullah et al. (2021); Sekhari et al. (2021).
Challenges of MU for LLMs
LLM unlearning introduces new challenges and complexities. First, LLMs are trained on massive amounts of data, which can unintentionally introduce biases and the memorization of personal and confidential information. Accordingly, it becomes challenging to precisely define and localize the ‘unlearning targets’, such as the subset of the training set or a knowledge concept that needs to be removed. Therefore, current studies on LLM unlearning Lu et al. (2022); Jang et al. (2022); Ilharco et al. (2022); Eldan & Russinovich (2023); Wu et al. (2023b); Yu et al. (2023); Zhang et al. (2023c); Yao et al. (2023) are typically context and task-dependent. There is a lack of standardized corpora for LLM unlearning. Second, the growing size of LLMs and the rise of black-box access to LLM-as-a-service present challenges in developing scalable and adaptable MU techniques to LLMs Bucknall & Trager (2023); Casper et al. (2024a). This also affects performance evaluation, given the absence of retraining as a benchmark. To address these challenges, previous studies have proposed approaches like in-context unlearning Pawelczyk et al. (2023) and fictitious unlearning Maini et al. (2024), where the former enables unlearning on black-box models, and the latter provides a synthetic for ease of retraining. Third, the scope of unlearning is often underspecified for LLMs. This issue is similar to challenges faced in model editing Mitchell et al. (2022). For instance, effective unlearning should ensure that LLMs delete knowledge of the targeted data within the predefined scope while simultaneously maintaining its utility for data outside of this scope. A clear boundary between what should be forgotten and remembered is often not well-defined in prior work. Fourth, despite the potential of LLM unlearning in diverse applications, there is a notable absence of comprehensive and reliable evaluation. For example, recent studies Shi et al. (2023); Patil et al. (2024); Lynch et al. (2024); Zhang et al. (2024a) have demonstrated that sensitive information can be reverse-engineered from an LLM post-unlearning, even if efforts were made to delete this information. This highlights the need for thorough and adversarial evaluations and the design of more mechanistic methods to guarantee the authenticity of unlearning.
@@ 3 Unpacking LLM Unlearning
In light of the existing literature on unlearning (Bourtoule et al., 2021; Jia et al., 2023; Kurmanji et al., 2023), and its progression in LLMs (Pawelczyk et al., 2023; Yao et al., 2023; Ishibashi & Shimodaira, 2023; Maini et al., 2024; Li et al., 2024a), we define the problem of LLM unlearning below.
@@
TOFU: Unlearning biographical
knowledge about fictitious authors
@@
(Thaker et al., 2024)
Harry Potter questions and
author biography in TOFU (Maini et al., 2024)
Guardrailing with a separate LLM
(I) Q&A about Harry Potter and unlearning authors
(O) Standard NLP benchmarks
N/A
(Zhang et al., 2024b)
Fictitious unlearning using TOFU (Maini et al., 2024)
Negative preference optimization
Same as TOFU (Maini et al., 2024)
N/A
(Li et al., 2024a)
Hazardous knowledge in the domain of
biology, cybersecurity, and chemistry
Optimization towards random
representations for unlearning concept
(I) Zero-shot Q&A about hazardous knowledge
(O) Zero-shot Q&A about other general knowledge,
and fluency of models
N/A
(Barbulescu & Triantafillou, 2024)
Specific text sequences memorized by LLM
Memorization-aware gradient ascent
(I) Memorization scores of the forget samples
(O) Commonsense and scientific reasoning tasks
N/A
(Wang et al., 2024c)
Private, toxic, and copyrighted knowledge
Factual relation removal in MLP layers
(I) Accuracy of generating ground-truth knowledge
(O) Evaluation on reasoning abilities
N/A
(Wang et al., 2024a)
Fictitious unlearning using TOFU (Maini et al., 2024)
Reverse KL divergence based
knowledge distillation
(I) Q&A about the unlearning authors
(O) Commonsense and scientific reasoning tasks
N/A
(Liu et al., 2024)
Fictitious unlearning using TOFU (Maini et al., 2024),
hazardous knowledge using WMDP (Li et al., 2024a),
copyrighted content in news articles and book
Detecting the forget prompts and
corrupting their embedding space
(I) Q&A or completion of the unlearned knowledge
(O) Eleven common LLM benchmarks
Runtime cost
@@
(1) Unlearning targets: Unlearning tasks may take on various forms and are closely related to the unlearning objectives. For instance, one could focus on data influence removal, while the other could emphasize model capability removal. Although these two aspects are intertwined, the former is often crucial for intellectual property (IP) protection, while the latter is more practical for AI alignment and safety. The literature identifies unlearning targets as specific data points, which could involve content containing harmful, unethical, or illegal language (Jang et al., 2022; Wu et al., 2023b). They have also been represented by higher-level unlearned knowledge, expressed through an unwanted text prompt or concept (Lu et al., 2022; Yao et al., 2023; Eldan & Russinovich, 2023). For example, the existing work (Eldan & Russinovich, 2023) defined the unlearning target as ‘Harry Potter’-related content, with the objective to avoid generating such content irrespective of where the content was learned: from the copyrighted material, blog posts, or news articles.
@@
(2) Influence erasure: Erasing the influence of unlearning targets and associated model capabilities requires a joint examination of both data and model influences rather than studies in isolation. Specifically, it is important to scrutinize the contributions of data sources to undesired model outputs, as well as the roles played by individual components within a model in generating these undesirable outcomes. This dual examination allows us to gain a more comprehensive understanding of the mechanisms driving these outputs, thereby facilitating the development of unlearning strategies to prevent them effectively. The objective of achieving complete influence erasure also implies the importance of robustness and generalization in unlearned behavior. When evaluating LLM unlearning, especially when using approximate methods shown in Table 1, a rigorous criterion is needed. Recent studies (Patil et al., 2024; Lynch et al., 2024) have underscored this viewpoint by demonstrating that forgotten information can be regenerated from LLMs post-unlearning using extraction or jailbreaking attacks.
(3) Unlearning effectiveness: The effectiveness of LLM unlearning extends beyond merely diminishing the influence of specific data points. A crucial aspect of effectiveness is the unlearning scope, as inspired by the editing scope (Mitchell et al., 2022). The unlearning scope defines the accuracy of influence erasure for in-scope examples, as well as the generation consistency for out-of-scope examples. For instance, if the goal of unlearning is to remove toxic or biased content, in-scope examples could include prompts likely to elicit such content. Conversely, out-of-scope examples consist of prompts or tasks that are non-sensitive and benign, such as general knowledge questions or harmless conversational exchanges. Differentiating between in-scope and out-of-scope examples for unlearning is often a difficult problem, as it requires determining when facts logically imply one another (Hase et al., 2023b; Cohen et al., 2023). This is also known as knowledge entanglement (Maini et al., 2024), where the unlearning targets and non-targets are closely related. Some methods have been shown to struggle to resolve such entanglement in such settings (Maini et al., 2024; Li et al., 2024a). In Table 1 (‘Effectiveness’ column), we have summarized the in-scope and out-of-scope examples from existing unlearning tasks in the literature.
@@
(4) Unlearning efficiency & feasibility: The majority of current research efforts have focused on developing rapid unlearning methods for LLMs due to the significant re-training costs involved (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023). Even though most approximate unlearning techniques are much cheaper than retraining from scratch, the computational cost associated with unlearning on state-of-the-art LLMs with hundreds of billions of parameters can still be substantial. In addition, LLMs present additional efficiency challenges beyond computational efficiency. These include the complexity and, at times, the infeasibility of pinpointing and attributing training data points designated for unlearning. Additionally, there is the challenge of executing unlearning in the context of black-box LLMs (Pawelczyk et al., 2023), where interactions with models are constrained to input-output queries.
@@
Mathematical modeling
Building upon the high-level LLM unlearning formulation presented earlier, we next provide mathematical modeling details and discuss the associated design choices. To facilitate comprehension, we provide a commonly-used formulation of LLM unlearning problems below. While this may not be the sole or optimal approach to LLM unlearning, it incorporates several key elements that are essential to the problem setup.
@@
In the dataset setup of LLM unlearning, we typically assume access to a forget set ($\mathcal{D}_{\mathrm{f}}$), the influence of which should be eliminated in LLM generation. For instance, $\mathcal{D}_{\mathrm{f}}$ might consist of a collection of harmful or toxic prompt-response pairs designated for degeneration (Yao et al., 2023). Moreover, if the original training set is available, then $\mathcal{D}_{\mathrm{f}}$ can be composed of a subset of training data points related to the unlearning target. Alternatively, it can be generated using synthesized data points based on a higher-level unlearned knowledge concept, or it can be derived from a set of extracted training data points reverse-engineered from the given LLM itself. In practice, the forget set $\mathcal{D}_{\mathrm{f}}$ is not required to belong precisely to the LLM’s training corpus. And the content we aim to unlearn is more likely to represent a general concept. Thus, LLM unlearning needs to not only unlearn specific training samples but also generalize to similar samples that share common characteristics.
@@
Besides the forget set $\mathcal{D}_{\mathrm{f}}$, there is usually a need for a retain set ($\mathcal{D}_{\mathrm{r}}$), which contains samples that are not subject to unlearning and used to preserve the utility of the unlearned model. Through the lens of the unlearning scope we discussed earlier, the forget set ($\mathcal{D}_{\mathrm{f}}$) provides in-scope examples earmarked for unlearning, while the retain set ($\mathcal{D}_{\mathrm{r}}$) involves examples out of the unlearning scope. Some recent studies have also attempted to develop LLM unlearning approaches that operate independently of access to forget and/or retain sets (Pawelczyk et al., 2023; Li et al., 2023b).
@@
We next introduce the model and optimization setups for LLM unlearning. Unlearning is often performed at the post-model training phase. As shown in (2), a common unlearning objective is to efficiently update the original pre-trained model so that the updated model can unlearn on $\mathcal{D}_{\mathrm{f}}$ while retaining its generation capability on $\mathcal{D}_{\mathrm{r}}$. Regarding the choice of optimizer to solve problem (2), recent work (Jia et al., 2024) has shown that using second-order optimization, such as Sophia (Liu et al., 2023a), yields better unlearning performance compared to first-order optimization. In addition, another design element is unlearning response ($y_{\mathrm{f}}$), referred to as the response of an unlearned model to in-scope examples. For example, in the stateful LLM unlearning method aimed at erasing information related to ‘Who’s Harry Potter?’ (Eldan & Russinovich, 2023), the unlearning response is based on word replacements using generic translations, like substituting ‘Quidditch’ with ‘Skyball’, as part of the unlearning process. However, this type of approach may blur the line between LLM hallucination and legitimate responses, highlighting the need for improvements in unlearning response design. Another choice is to specify $y_{\mathrm{f}}$ as empty response (Wu et al., 2023b; Patil et al., 2024), given by the rejection ‘I don’t know’ (Patil et al., 2024) or the customized response by ‘masking’ the unlearning information in (Wu et al., 2023b). However, we need to ensure that the empty response targets only examples within the unlearning scope. Otherwise, frequent rejections may occur, potentially diminishing the user experience with LLMs.
4 Current Unlearning Techniques and Overlooked Principles